OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sqrt{x^2-x+1}-ax-b\] , a and b are positive constants

OpenStudy (anonymous):

@aaronq ?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

you can break this up...

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\sqrt{x^2-x+1}-ax-b=\lim_{x \rightarrow \infty}\sqrt{x^2-x+1}-\lim_{x \rightarrow \infty}ax-\lim_{x \rightarrow \infty}b\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well, that may not do what you need... let me think for a sec.

ganeshie8 (ganeshie8):

rationalizing the numerator may help

ganeshie8 (ganeshie8):

\[\lim_{x \rightarrow \infty} \dfrac{ x^2-x+1-(ax+b)^2 }{ \sqrt{x^2-x+1}+ax+b} \]

OpenStudy (anonymous):

you could do this... \[\sqrt{x^2-x+1}-(ax+b)=\sqrt{x^2-x+1}-(ax+b)\cdot\frac{ \sqrt{x^2-x+1}+(ax+b) }{ \sqrt{x^2-x+1}+(ax+b) }\]

ganeshie8 (ganeshie8):

apply L'hospitals now

OpenStudy (anonymous):

on your equation @ganeshie8 ?

ganeshie8 (ganeshie8):

its getting complicated... one sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

u der? @ganeshie8 ?

OpenStudy (anonymous):

@pgpilot326 ?

OpenStudy (anonymous):

multiply out and gather like terms for the numerator. then split into separate limits? i think this might do it.

ganeshie8 (ganeshie8):

\[\lim_{x \rightarrow \infty} \dfrac{ x^2(1-a)+x(-1-2ab)+1-2ab }{ \sqrt{x^2-x+1}+ax+b} \] the limit is DNE unless \(1-a = 0\)

ganeshie8 (ganeshie8):

when a = 1, the limit will be \(\dfrac{1}{2}-b\) for any other value of a, the limit DNE

OpenStudy (anonymous):

how did u get it?

ganeshie8 (ganeshie8):

for \(a \ne 1\), numerator's degree is larger than denominator, so as x increases, the expression grows without any bound... consequently the limit wont exist

OpenStudy (anonymous):

u applied LH?

OpenStudy (anonymous):

\[\sqrt{x^2-x+1}<x,\,\, \forall x>1\]

ganeshie8 (ganeshie8):

^^ plugin a=1, then apply LH or divide numerator and denominator by x... anything will do

OpenStudy (anonymous):

what did u do? and how can we plug a=1?

ganeshie8 (ganeshie8):

\[\lim_{x \rightarrow \infty} \dfrac{ x^2(1-a)+x(-1-2ab)+1-2ab }{ \sqrt{x^2-x+1}+ax+b}\] we consider two cases : \(a \ne 1\) \(a = 1\)

OpenStudy (anonymous):

k

ganeshie8 (ganeshie8):

for \(a \ne 1\) case, clearly the numerator's degree is 2, and the denominator's degree is 1. so the limit is DNE. (is this convinsing ? )

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so the limit for \(a\ne 1\) case is DNE. lets work \(a = 1\) case

OpenStudy (anonymous):

ok nw if a=1 the degree will become same

ganeshie8 (ganeshie8):

\(a = 1\) case : \[\lim_{x \rightarrow \infty} \dfrac{ x(-1-2b)+1-2ab }{ \sqrt{x^2-x+1}+x+b} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got that the limit would exist if a=1 ! but what is the limit?

ganeshie8 (ganeshie8):

divide numerator and denominator by x

ganeshie8 (ganeshie8):

\[\lim_{x \rightarrow \infty} \dfrac{ (-1-2b)+(1-2ab)/x }{ \sqrt{1-1/x+1/x^2}+1+b/x} \]

ganeshie8 (ganeshie8):

take the limit ^^

OpenStudy (anonymous):

-1-2b/2

ganeshie8 (ganeshie8):

Yep !